I'm kind of confused how to solve by elimination. Can someone help me?
3x+4y=-1
-9x-4y=13

Answer by josgarithmetic(39617)

Original System
3x+4y=-1
-9x-4y=13

3(3x+4y)=3(-1)
-9x-4y=13
ADD.
Solve y.

3x+4y=-1
-9x-4y=13
ADD.
Solve x.